Circular No. 9127

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

CBAT Director: Daniel W. E. Green; Room 209; Dept. of Earth and Planetary Sciences; Harvard University; 20 Oxford St.; Cambridge, MA 02138, U.S.A.

cbat@iau.org  cbatiau@eps.harvard.edu  ISSN 0081-0304

                URL http://www.cfa.harvard.edu/iau/cbat.html

COMET C/2010 E5 (SCOTTI) J. V. Scotti, Lunar and Planetary Laboratory, reports his discovery of

a diffuse comet with a short tail extending toward the northwest on CCD mosaic images obtained with the 0.9-m f/3 Spacewatch reflector at Kitt Peak (discovery observation tabulated below). Nine co-added exposures by Scotti on Mar. 14.3 UT with the Spacewatch 1.8-m reflector show a faint tail clearly visible extending 0′.62 in p.a. 304^{o} and a small, slightly diffuse coma of diameter 8′′.4; stacked images on Mar. 15 show a coma diameter of 10′′.3 and a tail 1′.09 long in p.a. 306^{o} (the end of the tail in a lesscrowded star region than on Mar. 14). T. H. Bressi finds that images with the same telescope on Mar. 17.4 show a tail in p.a. 320^{o}. After posting on the Minor Planet Center’s ‘NEOCP’ webpage, numerous other CCD astrometrists have commented on the object’s cometary appearance. P. Holvorcem, Campinas, Brazil, writes that three 270-s co-added exposures taken with the Tenagra 0.81-m telescope around Mar. 15.36 show a coma ∼ 7′′ in diameter and a narrow, straight, 17′′ tail toward p.a. 291^{o}, while three co-added 360-s exposures taken around Mar. 17.34 under poorer seeing show a coma of diameter ∼ 9′′ and a 5′′ tail toward p.a. 280^{o}. L. Buzzi (Varese, Italy, 0.60-m reflector, Mar. 17.0–17.1) notes that stacked images reveal a 10′′ condensed coma and a tail at least 25′′ long in p.a. 305^{o}. S. Foglia reports that 18 stacked 60-s images taken by R. Holmes (Ashmore, IL, U.S.A., 0.61-m reflector, Mar. 17.24–17.29) show an 8′′ coma and a 40′′ tail in p.a. 315^{o}. R-band images taken with the Magdalena Ridge 2.4-m reflector by W. H. Ryan and E. V. Ryan on Mar. 17.4 show a distinct tail at p.a. ∼ 300^{o}. L. Elenin finds that six stacked 300-s unfiltered CCD exposures taken with a 0.36-m astrograph at Tzec Maun Observatory (near Mayhill, NM, USA) on Mar. 17.44 show a 4′′ diffuse coma and a 3′′.5 tail in p.a. 315^{o}. Twenty 120-s unfiltered exposures taken remotely by E. Guido and G. Sostero with a 0.25-m reflector near Mayhill on Mar. 17.5 show a coma nearly 10′′ in diameter with a broad tail ∼ 25′′ long toward the northwest.

2010 UT α_{2000} δ_{2000} Mag. Observer
Mar. 14.23418 12^{h}40^{m}39.^{s}80 +1^{o}31′55′′.1 20.1 Scotti

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2010-F14.

T = 2011 Jan. 20.812 TT  ω = 218.^{o}367
Ω = 13.857 }2000.0
q = 4.02520 AU  i = 27.867
2010 March 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2010 CBAT  Daniel W. E. Green